OpenStudy (anonymous):

2cos(x) is less than or equal to cos(x-pi/2) solve for [0,2pi) show the algebraic work to test the intervals

OpenStudy (anonymous):

\[2cosx \le \cos(x-\Pi/2)\] \[0 \le x < 2\Pi\]

OpenStudy (anonymous):